# Lite-Mind: Towards Efficient and Robust Brain Representation Learning

### v3

@@ 1. Introduction
The success of fMRI-to-image retrieval heavily relies on aligning fMRI signals with the representation space of images. However, fMRI data often suffers from spatial redundancy, noise, and sample sparsity, leading to poor representation of fMRI signals and potential overfitting to noise distribution for deep models. To address these challenges, previous studies have found ridge regression and shallow linear models to be effective solutions for fMRI-based models (Pereira et al., 2018; Kewenig et al., 2023; Emmanuele et al., 2021; Allen et al., 2022; Horikawa and Kamitani, 2017; Zou et al., 2022). Recently, it has been demonstrated the effectiveness of leveraging pretrained CLIP models (Radford et al., 2021) as a powerful bridge between fMRI voxels and images, since CLIP’s image embeddings capture fine-grained and semantic information of pictures (Liu et al., 2023; Mai and Zhang, 2023; Lin et al., 2022; Scotti et al., 2023). Initially, the CLS embeddings of the CLIP image encoder was used to align voxels in Mind-Reader (Lin et al., 2022). Considering the CLS embeddings alone provide limited information for fMRI data, BrainClip (Liu et al., 2023) and UniBrain (Mai and Zhang, 2023) introduce the text annotations of images to assist in aligning fMRI voxels with image representations. Brain-diffuser (Ozcelik and VanRullen, 2023) aligns voxel representations with the last hidden layer of CLIP using large-scale linear models ($257\times 14$M) for the first time.
@@ 2.2. Fourier Transform in Deep Learning
Fourier Transform plays a vital role in the area of digital signal processing. It has been introduced to deep learning for enhanced learning performance (Gueguen et al., 2018; Ehrlich and Davis, 2019; Yang and Soatto, 2020), and has demonstrated excellent performance in many fields, such as computer vision, natural language processing, and time series analysis. In the field of computer vision, GFNet (Rao et al., 2021) utilizes Fast Fourier Transform(FFT) to convert images to the frequency domain and exchange global information between learnable filters. As a continuous global convolution independent of input resolution, Guibas et al. (Guibas et al., 2021) design the Adaptive Fourier Neural Operator(AFNO) frame token mixing. Xu et al. (Xu et al., 2020) devise a learning-based frequency selection method to identify trivial frequency components and improve image classification accuracy. As for natural language processing, Lee-Thorp et al. (Lee-Thorp et al., 2021) use the Fourier Transform as a text token mixing mechanism for text classification. In research of time series forecast, Fourier Transform has deeper applications  (Cao et al., 2020; Lange et al., 2021; Koç and Koç, 2022; Yang and Hong, 2022). To increase the accuracy of multivariate time-series forecasting, Cao et al. (Cao et al., 2020) propose a spectral temporal graph neural network (StemGNN), which mines the correlations and time dependencies between sequences in the frequency domain. Yang et al. (Yang and Hong, 2022) propose bilinear temporal spectral fusion (BTSF), which updates the feature representation in a fused manner by explicitly encoding time-frequency pairs and using two aggregation modules: spectrum-to-time and time-to-spectrum. Yi et al. (Yi et al., 2024) prove that frequency-domain MLP is more efficient than time-domain MLP. Processing fMRI frequency domain information is more helpful for understanding the mechanism of collaborative operation within the human brain. However, there is currently no research on processing fMRI signals in the frequency domain for representation learning.
@@ 5.4. Ablations and visualization
Embeddings Dimension Parameters Retrieval
[…]
ViT-B/32 CLS 512 6.7M 57.5% 61.8%
ViT-L/14 CLS 768 6.7M 60.3% 64.2%
ViT-B/32 Hidden 50×512 10.4M 91.1% 96.5%
ViT-L/14 Hidden 257×768 12.5M 94.6% 97.4%
@@
6. Limitations and Future Work
Due to inherent differences between Subjects, Lite-Mind still needs to train Subject-specific models. Although our exploratory experiments have shown that a common model can achieve comparable retrieval accuracy on four Subjects, it’s still necessary to ensure sufficient training data. Further design beyond the model level is still needed for cross-Subject models, to achieve Subject-specific few-shot or even zero-shot. In addition, Lite-Mind performs worse on Subjects with shorter voxel lengths than on Subjects with longer, consistent with previous research, which lacks relevant neuroscience theoretical evidence. This provides a promising avenue for the interpretability of deep learning in neuroscience.
@@ 7. Conclusion
In this work, we propose Lite-Mind, an extremely lightweight brain representation learning paradigm based on Fourier Transform for cross-modal fMRI-to-image retrieval. Our DFT Backbone is an efficient means of obtaining fine-grained representation alignment between fMRI signals and visual stimuli. With high retrieval accuracy, Lite-Mind can transform many downstream fMRI tasks into retrieval-based tasks, such as zero-shot classification. Meanwhile, Lite-Mind’s dependence on dataset size and voxel length is less than larger models, demonstrating excellent robustness across different fMRI datasets.

### v4

@@ Abstract.
fMRI, Brain-computer Interface (BCI), Cross-modal Retrieval
copyright: acmlicensedjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681229isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
The success of fMRI-to-image retrieval heavily relies on aligning fMRI signals with the representation space of images, following a series of cross-modal alignment efforts (Ramesh et al., 2022; Wu et al., 2024; Radford et al., 2021). However, fMRI data often suffers from spatial redundancy, noise, and sample sparsity, leading to poor representation of fMRI signals and potential overfitting to noise distribution for deep models. To address these challenges, previous studies have found ridge regression and shallow linear models to be effective solutions for fMRI-based models (Pereira et al., 2018; Kewenig et al., 2023; Emmanuele et al., 2021; Allen et al., 2022; Horikawa and Kamitani, 2017; Zou et al., 2022). Recently, it has been demonstrated the effectiveness of leveraging pretrained CLIP models (Radford et al., 2021) as a powerful bridge between fMRI voxels and images, since CLIP’s image embeddings capture fine-grained and semantic information of pictures (Liu et al., 2023; Mai and Zhang, 2023; Lin et al., 2022; Scotti et al., 2023; Zhang et al., 2024). Initially, the CLS embeddings of the CLIP image encoder was used to align voxels in Mind-Reader (Lin et al., 2022). BrainClip (Liu et al., 2023) and UniBrain (Mai and Zhang, 2023) introduce the text annotations of images to assist in aligning fMRI voxels with image representations. Brain-diffuser (Ozcelik and VanRullen, 2023) aligns voxel representations with the last hidden layer of CLIP using large-scale linear models ($257\times 14$M) for the first time.
@@ 2.2. Fourier Transform in Deep Learning
Fourier Transform plays a vital role in the area of digital signal processing. It has been introduced to deep learning for enhanced learning performance (Yi et al., 2023a; Lao et al., 2024; Yang and Soatto, 2020; Zhang et al., 2024), and has demonstrated excellent performance in many fields, such as computer vision, natural language processing, and time series analysis. In the field of computer vision, GFNet (Rao et al., 2021) utilizes Fast Fourier Transform (FFT) to convert images to the frequency domain and exchange global information between learnable filters. As a continuous global convolution independent of input resolution, Guibas et al. (Guibas et al., 2021) design the Adaptive Fourier Neural Operator (AFNO) frame token mixing. Xu et al. (Xu et al., 2020) devise a learning-based frequency selection method to identify trivial frequency components and improve image classification accuracy. As for natural language processing, Lee-Thorp et al. (Lee-Thorp et al., 2021) use the Fourier Transform as a text token mixing mechanism for text classification. In research of time series forecast, Fourier Transform has deeper applications  (Yi et al., 2023b; Lange et al., 2021; Koç and Koç, 2022). To increase the accuracy of multivariate time-series forecasting, Cao et al. (Cao et al., 2020) propose a spectral temporal graph neural network (StemGNN), which mines the correlations and time dependencies between sequences in the frequency domain. Yang et al. (Yang and Hong, 2022) propose bilinear temporal spectral fusion (BTSF), which updates the feature representation in a fused manner by explicitly encoding time-frequency pairs and using two aggregation modules: spectrum-to-time and time-to-spectrum. Yi et al. (Yi et al., 2024) prove that frequency-domain MLP is more efficient than time-domain MLP. Processing fMRI frequency domain information is more helpful for understanding the mechanism of collaborative operation within the human brain. However, there is currently no research on processing fMRI signals in the frequency domain for representation learning.
@@ 5.4. Ablations and visualization
6. Limitations
Due to inherent differences between Subjects, Lite-Mind still needs to train Subject-specific models. Although our exploratory experiments have shown that a common model can achieve comparable retrieval accuracy on four Subjects, it’s still necessary to ensure sufficient training data.
@@ 7. Conclusion
In this work, we propose Lite-Mind, an extremely lightweight brain representation learning paradigm based on Fourier Transform for fMRI-to-image retrieval. Our DFT Backbone is an efficient means of obtaining fine-grained representation alignment between fMRI signals and visual stimuli. With high retrieval accuracy, Lite-Mind can transform downstream fMRI tasks into retrieval-based tasks, such as zero-shot classification. Meanwhile, Lite-Mind’s dependence on dataset size and voxel length is less than that of larger models, demonstrating excellent robustness across different fMRI datasets.
8. Acknowledgments
This work is supported by the National Key Research and Development Program of China (No. 2022YFB3104700), the National Natural Science Foundation of China (No. 61976158, No. 62376198), Shanghai Baiyulan Pujiang Project (No. 08002360429).
Appendix A Additional Method Details
A.1. Implementation details
Data preprocessing. We downloaded the NSD dataset from the official website and used Takagi’s code to extract nsdgenal fMRI voxels, while Takagi extracted the stream region of the NSD dataset. We noticed that MindEye scaled the fMRI voxel values in advance, while we did not. The difference in fMRI voxels input data is shown in the example in Figure 6. NSD image files come from nsd-stimuli.hdf5 file and have a unified size of $425\times 425$. We did not perform any data augmentation on the image and straightly extracted the hidden layer representation (size of $257\times 768$) of the image through CLIP ViT-L/14 for training.
Hyper-parameters. On the NSD dataset, during training DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, $\alpha=1$, and CLIP’s contrastive loss is unidirectional for image retrieval and fMRI retrieval. Owning to the lightweight nature of Lite-Mind, the batch size is set to 500, the learning rate is 1e-3, patch size is set to 480, and Filter library size is 4. Filter block layers are the same for all subjects.
About LAION-5B retrieval, for DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, while the weight decay is 6.02e-2 for diffusion projector. CLIP’s contrastive loss is unidirectional for image retrieval, $\alpha=0.5$, the batch size is set to 80 and the learning rate is 1.16e-3 for DFT Backbone while 1.1e-4 for diffusion projector. Note that hyper-parameters in Diffusion Projector are the value of the open-source DALLE·2 from https://github.com/lucidrains/DALLE2-pytorch.
On the GOD dataset, the hyperparameters are the same as those on the NSD dataset, except for the batch size and patch size which are set to 1200 and 8 for all 5 subjects on the GOD dataset, respectively.
(a) MindEye
(b) Lite-Mind
Figure 6. The fMRI averaged activities of MindEye and Lite-Mind responding to the same image, respectively. The figures only visualize the activities of the first 1000 voxels for illustration.
Appendix B Additional Experiment Results
B.1. Additional Results of Retrieval
To demonstrate the applicability of our method, we also conducted experiments on three other Subjects, i.e., Subject 2, Subject 5, and Subject 7, on the NSD dataset, and the experimental results are shown in Table 7. In order to control model size similarity, we did not make more targeted model adjustments for subjects with shorter voxel lengths. It can be seen that the retrieval accuracies of Subjects 1 and 2 are greater than 90%, and the retrieval accuracies of Subjects 5 and 7 are also above 80%. The results prove that DFT Backbone can efficiently work on different subjects. Note that MindEye only includes the results of the overall model in Subjects 2, 5, and 7 in its Appendix, and it does not evaluate the effect of individual MLP Backbone. As a result, Table 7 does not present any evaluation results for MindEye.
Method Voxel Length Parameters Retrieval
[…]
Lite-Mind(Subj 1) 15724 12.51M 94.6% 97.4%
Lite-Mind(Subj 2) 14278 12.49M 94.1% 98.2%
Lite-Mind(Subj 5) 13039 12.47M 80.5% 86.3%
Lite-Mind(Subj 7) 12682 12.47M 81.7% 82.3%
Table 7. Addtional retrieval performance for individual subjects on 982 test images of the NSD dataset.
Method Low-Level High-Level
[…]
Lite-Mind(Subject 1) .134 .332 78.8% 88.9% 88.5% 88.8% .730 .451
Lite-Mind(Subject 2) .120 .328 78.0% 89.4% 86.3% 87.4% .730 .446
Lite-Mind(Subject 5) .123 .332 79.4% 90.0% 88.8% 89.9% .712 .440
Lite-Mind(Subject 7) .121 .331 78.7% 88.8% 87.8% 88.5% .723 .448
Table 8. LAION-5B retrieval alternative reconstruction performance for the specific subject.
B.2. Additional Results of LAION-5B Retrieval
In order to better reflect the retrieval performance of Lite-Mind on LAION-5B, we presented the performance indicators of LAION-5B retrieval substitution reconstruction for other subjects, i.e., Subject 2, Subject 5, and Subject 7, in Table 8, corresponding to the average performance of subjects in Table 1. Similarly, the visualization results of the other subjects in Figure 7 correspond to the image samples of Subject 1 in Figure 4. Based on the comprehensive table and graph, it can be found that Lite-Mind has good generalization on all four subjects, verifying the LAION-5B retrieval ability of Lite-Mind on different subjects. Meanwhile, as shown in Figure 7, the retrieval performance of LAION-5B completely depends on the retrieval accuracy of the CLS model in the test set, such that images retrieved incorrectly in the test set may also have retrieval bias on LAION-5B, for example treating a teddy bear as an image of a cat or dog as shown in Figure 7. However, both MindEye and Lite-Mind exhibit relatively low retrieval accuracy with aligned CLS embedding models. In the future, it would be beneficial to explore models that improve the alignment of CLS embeddings or employ more efficient methods to directly perform retrieval through hidden layers in LAION-5B.
Figure 7. Addtional retrieval results corresponding to Figure 4. The left column marked by a red box in every two columns represents the original image seen by the subject, and the right column represents the image retrieved on LAION-5B.
B.3. Additional Results of Zero-shot Classification
In the main body of the paper, we only demonstrated the zero-shot classification effect of Lite-Mind on the GOD dataset. The corresponding retrieval results of each Subject are shown in the Table 9 below.
[…]
top1 top5
Lite-Mind(Subj 1) 4466 15.50M 30.0% 60.0%
Lite-Mind(Subj 2) 4404 15.46M 38.0% 58.0%
Lite-Mind(Subj 3) 4643 15.64M 38.0% 72.0%
Lite-Mind(Subj 4) 4133 15.24M 42.0% 62.0%
Lite-Mind(Subj 5) 4370 15.43M 26.0% 54.0%
Table 9. Retrieval performance on the GOD dataset.
B.4. Hyper-parameters Experiment
In this section, we explore the influence of some hyperparameters on the model, including patch size and number of filters $M$, to verify the parameter sensitivity of the model. All the experimental results on the NSD dataset are from Subject 1, with a retrieving pool size of 300.
Patch size. We conducted experiments by varying the patch size as presented in Table 10. The results exhibit stable retrieval accuracy above 90%, which indicates that our DFT Backbone is not sensitive to part size. Since we found patch size has less effect on retrieval accuracy, we chose a relatively large patch size (i.e., 480) to ensure fewer parameters and faster convergence.
Patch size Parameters Retrieval
[…]
50 14.0M 91.6% 96.4%
200 12.7M 92.9% 96.9%
480 12.5M 94.6% 97.1%
600 12.3M 94.2% 97.3%
900 11.9M 92.1% 95.5%
Table 10. Retrieval performance for different patch size on the NSD dataset.
Filter library size. We conducted experiments by varying the Filter library size as presented in Table 11. Specifically, there is a significant performance improvement from $M=1$ to $M=4$, while a slight fluctuation is observed for $M=4/8$. By setting $M=4$, the model has the ability to acquire diverse and distinct feature patterns from various dimensions of the frequency response while still maintaining an appropriate computational cost. Therefore, we determine that $M=4$ is the optimal choice on the NSD dataset.
Filters Parameters Retrieval
[…]
1 9.0M 89.8% 96.4%
2 10.2M 93.5% 97.4%
4 12.5M 94.6% 97.1%
8 17.2M 94.4% 96.7%
Table 11. Retrieval performance for different Filter library size on the NSD dataset.
Embeddings Dimension Parameters Retrieval
[…]
ViT-B/32 CLS 512 6.7M 57.5% 61.8%
ViT-L/14 CLS 768 6.7M 60.3% 64.2%
ViT-B/32 Hidden 50×512 10.4M 91.1% 96.5%
ViT-L/14 Hidden 257×768 12.5M 94.6% 97.4%
Figure 8. Lite-Mind’s training loss curve and testing accuracy curve for Suject 1, 2, 5, 7 on the NSD dataset. The testing accuracy is calculated based on a retrieval pool of size 982.
Figure 9. Lite-Mind’s retrieval heatmap on the NSD dataset for Subject 1. The larger figure on the right represents 982 × 982’s heatmap, and the smaller figure on the left represents the 50 × 50 subgraph.
Appendix C Additional Visualization
C.1. Training Curve
We visualized the training process of all four subjects on the NSD dataset in Figure 8. As the training epochs increased, the loss of the training set rapidly decreased, while the accuracy of the test set rapidly increased and then showed a slow upward trend. Accuracy refers to the hit rate of correct retrieval from 982 test set images.
C.2. Retrieval Heatmap
We visualized the retrieval heatmap for Subject 1 on all 982 test images of the NSD dataset in Figure 9. It can be observed that the similarity is highest on the diagonal, and the color of the retrieved heat map is darker. It shows that Lite-Mind has effectively retrieved corresponding images, even if there are many similar images in the test set, which verifies the fine-grained ability of Lite-Mind.
C.3. Information Alignment
We visualized the T-SNE plot between the voxel embeddings output by DFT Backbone and the image embeddings of frozen CLIP as the accuracy of the test set improved, as shown in Figure 13. We can observe that as the training progresses, the retrieval accuracy of the test set improves, and the shape of voxel embeddings tends to be closer to image embeddings, indicating the success of contrastive learning.
C.4. More Retrieval Cases
In this section, we visualize retrieval failure cases in Figure 10, although only dozens of images are not in the Top 1. From the left half of the Figure, it can be seen that though not in Top 1, Lite-Mind still retrieved ground-truth in Top 2, and images in Top 4 are similar(either semantically similar or structurally similar, eg. animals in the wild or an airplane on the runway). On the contrary, a smaller proportion of ground-truth did not appear within the Top 4, as shown in the right half of the Figure. From the perspective of the images themselves, most of the scenes are too complex, even abstract (as shown in the second image), which may be the reason for the retrieval failure.
Figure 10. Partial failure retrieval results of Lite-Mind on all 982 test images for Subject 1. The number below each image represents the similarity score.
C.5. Visualization in Frequency Domain
We visualized the weights of Filter library of Filter Blocks, as shown in Figure 11 and 12. Visualization is divided into the real part (left) and imaginary part (right) of filter weights. It can be observed that different filters have varying degrees of attention to different tokens, and the frequency domain better captures this characteristic. Interestingly, the weight of the imaginary part for the first and last tokens is almost always 0, indicating that the noise is distributed in these two tokens.
(a) The real part and imaginary of a filter weights.
(b) The real part and imaginary of another filter weights from the same Filter Library.
Figure 11. Weights visualization for DFT Backbone of $257\times 768$ embedding length.
(a) The real part and imaginary of a filter weights.
(b) The real part and imaginary of another filter weights from the same Filter Library.
Figure 12. Weights visualization for DFT Backbone of 768 CLS embedding length.
Appendix D Theoretical Analysis.
D.1. Complex Multiplication.
For two complex number values $\mathcal{Z}_{1}=(a+jb)$ and $\mathcal{Z}_{2}=(c+jd)$, where $a$ and $c$ is the real part of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ respectively, $b$ and $d$ is the imaginary part of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ respectively. Then the multiplication of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ is calculated by:
(10) $\mathcal{Z}_{1}\mathcal{Z}_{2}$ $=(a+jb)(c+jd)$
$=ac+j^{2}bd+jad+jbc$
$=(ac-bd)+j(ad+bc)$
D.2. Theorem Proof.
Theorem 1. Suppose that H is the representation of raw fMRI voxel tokens and $\mathcal{H}$ is the corresponding frequency components of the spectrum, then the energy of voxel tokens in the spatial domain is equal to the energy of its representation in the frequency domain. Formally, we can express this with the above notations:
(11) $\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v=\int_{-\infty}^{\infty}{|\mathcal{H}(f)|}^{2}\,\mathrm{d}f$
Where $\mathcal{H}(f)=\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v$, $v$ is the token dimension, $f$ is the frequency dimension.
Proof.
Given the representation of raw voxel token series $H\in\mathbb{R}^{C\times N}$, let us consider performing integration in the $N$ dimension (spatial dimension), denoted as the integral over $v$, then
(12) $\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v=\int_{-\infty}^{\infty}{\textbf{H}(v)}{\textbf{H}^{*}(v)}\,\mathrm{d}v$
where $\textbf{H}^{*}(v)$ is the conjugate of $\textbf{H}(v)$. According to IDFT, $\textbf{H}^{*}(v)=\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f$, we can obtain
(13) $\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v$ $=\int_{-\infty}^{\infty}{\textbf{H}(v)}[\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f]\,\mathrm{d}v$
$=\int_{-\infty}^{\infty}{\mathcal{H}}^{*}(f)[\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v]\,\mathrm{d}f$
$=\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}{\mathcal{H}(f)}\,\mathrm{d}f$
$=\int_{-\infty}^{\infty}{|\mathcal{H}(f)|}^{2}\,\mathrm{d}f$
Proved. ∎
Therefore, the energy of a voxel token series in the spatial domain is equal to the energy of its representation in the frequency domain.
Theorem 2. Given the voxel token series input H and its corresponding frequency domain conversion H, the operations of frequency-domain MLP on H can be represented as global convolutions on H in the spatial domain. This can be given by:
(14) $\mathcal{H}\mathcal{W}+\mathcal{B}=\mathcal{F}(\textbf{H}*W+B)$
Where $\mathcal{F}$ is DFT, * is a circular convolution, $\mathcal{W}$ and $\mathcal{B}$ are the complex number weight and bias, and $W$ and $B$ are the weight and bias in the spatial domain.
Proof.
Suppose that we conduct operations in the N (i.e., token dimension), then
(15) $\mathcal{F}(\textbf{H}(v)*W(v))=\int_{-\infty}^{\infty}{(\textbf{H}(v)*W(v))e^{-j2\pi fv}}\,\mathrm{d}v$
According to convolution theorem, $\textbf{H}(v)*W(v)=\int_{-\infty}^{\infty}(\textbf{H}(\tau)W(v-\tau))\,\mathrm{d}\tau$, then
(16) $\mathcal{F}(\textbf{H}(v)*W(v))$ $=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{(\textbf{H}(\tau)W(v-\tau))e^{-j2\pi fv}}\,\mathrm{d}\tau}\,\mathrm{d}v$
$=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(v-\tau)e^{-j2\pi fv}}\,\mathrm{d}v}\textbf{H}(\tau)\,\mathrm{d}\tau$
Let $x=v-\tau$,then
(17) $\mathcal{F}(\textbf{H}(v)*W(v))$ $=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(x)e^{-j2\pi f(x+\tau)}}\,\mathrm{d}x}\textbf{H}(\tau)\,\mathrm{d}\tau$
$=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(x)e^{-j2\pi fx}e^{-j2\pi f\tau}}\,\mathrm{d}x}\textbf{H}(\tau)\,\mathrm{d}\tau$
$=\int_{-\infty}^{\infty}{(\textbf{H}(\tau)*)e^{-j2\pi f\tau}}\,\mathrm{d}\tau\int_{-\infty}^{\infty}{(\textbf{W}(x)*)e^{-j2\pi fx}}\,\mathrm{d}x$
$=\mathcal{H}(f)\mathcal{W}(f)$
Accordingly, $\textbf{H}(v)*W(v)$ in the spatial domain is equal to $\mathcal{H}(f)\mathcal{W}(f)$ in the frequency domain. Therefore, the operations of FreMLP ($\mathcal{H}\mathcal{W}+\mathcal{B}$) in the token dimension (i.e., v = N ) are equal to the operations ($\mathcal{F}(\textbf{H}*W+B)$) in the spatial domain. This implies that frequency-domain MLPs can be viewed as global convolutions in the spatial domain. Proved. ∎
D.3. Complexity Analysis
For a fMRI voxel with a length of $l$, we divide it into $n$ patches. Assuming $L_{1}$ and $L_{2}$ are the layer depths of MLP Backbone and DFT Backbone respectively, the middle layer dimension of MLP Backbone is $D$, and the alignment embedding dimension is $n^{\prime}\times D^{\prime}$, where $n^{\prime}$ is the the number of tokens of CLIP. The time complexity of MLP Backbone is $O(lD+L_{1}D^{2}+n^{\prime}DD^{\prime})$. For DFT Backobone, the time complexity of patchify and tokenization is $O(lD^{\prime})$, and the time complexity of DFT, IDFT, and filtering for each layer is $O(2nD^{\prime}\operatorname{log}n+nD^{\prime})$. The time complexity of FreMLP is $O(2nD^{\prime}\operatorname{log}n+2nn^{\prime}D^{\prime}+2n^{\prime}D^{\prime})$. Thus the time complexity of the entire DFT Backbone is:
(18) $O(4nD^{\prime}\operatorname{log}n+(n+2nn^{\prime}+2n^{\prime})D^{\prime})$
$=O((n\operatorname{log}n+nn^{\prime}+n^{\prime})D^{\prime})$
Quantitative analysis algorithm complexity for DFT Backbone has been shown in Table 3.
Figure 13. T-SNE visualization between the voxel embeddings output by DFT Backbone and the image embedding of frozen CLIP. Accuracy in the title refers to the hit rate of correct retrieval from 982 test set images and the blue dots represent voxel embeddings.